OpenStudy (anonymous):

@amistre64 @zepdrix @paki I need help with Complex Fractions

OpenStudy (anonymous):

I need to write two complex fractions that simplify to \[\frac{ x-2 }{ x+4 }\]

OpenStudy (amistre64):

x or x^2 ?

OpenStudy (anonymous):

x

OpenStudy (amistre64):

multiply top and bottom by a complex number perhaps? or a conjugate ... but a conjugate tends to get us x^2

OpenStudy (unklerhaukus):

where \(f(x)\neq 0\)\[\frac{f(x)(x-2)}{f(x)(x+4)}= \frac{x-2}{x+4}\]

OpenStudy (amistre64):

yeh i was thinking along those lines :)

OpenStudy (anonymous):

Okay so that works. I don't necessarily need functions for this. I need any complex fraction that is similar to this one \[\frac{ \frac{ 2 }{ x } -\frac{ 3 }{ y } }{ \frac{ -5 }{ x } + \frac{ 7 }{ y } }\]

OpenStudy (amistre64):

complex, i was thinking you needed a complex number like 2+3i .... you simply want some complicated fractional working instead.

OpenStudy (anonymous):

Yeah unfortunately. That's how my textbook goes by

OpenStudy (amistre64):

divide top and bottom by say: xy, then split fractions ... an idea

OpenStudy (anonymous):

Okay

OpenStudy (amistre64):

\[\frac{ \frac{x-2}{xy} }{ \frac{x+4 }{xy}}\] \[\frac{ \frac{x}{xy}-\frac{2}{xy} }{ \frac{x}{xy}+\frac{4 }{xy}}\] \[\frac{ \frac{1}{y}-\frac{2}{xy} }{ \frac{1}{y}+\frac{4 }{xy}}\] etc ...

OpenStudy (unklerhaukus):

or multiply \[\frac{ \frac{ 2 }{ x } -\frac{ 3 }{ y } }{ \frac{ -5 }{ x } + \frac{ 7 }{ y } }= \frac{( \frac{ 2 }{ x } -\frac{ 3 }{ y } )\times xy}{ (\frac{ -5 }{ x } + \frac{ 7 }{ y } )\times xy}=\]

OpenStudy (anonymous):

Oh, okay. Now, I understand. Thank you both!

OpenStudy (amistre64):

good luck